OpenStudy (anonymous):

how do you find the magnitude of gravity if your only given the coefficients of static and kinetic friction and the weight of a box at rest?

OpenStudy (anonymous):

If you mean the magnitude of the gravitational force on the box, that's just the weight. However, I assume you mean you want to find the magnitude of the acceleration due to gravity, g ~ 9.8 m/s^2. Here's how you'd do it. If you cause the box to accelerate by applying a force, then \[F_{net} = ma \] but you have kinetic friction acting against you, so \[F_{net} = F_{applied} - \mu_k W = ma \] where W is the weight of the box. You can measure the acceleration -- This allows you to find the mass of the box: \[ m = \frac{F_{applied} - \mu_k W}{a} \] Last by not least, you know that the weight W = m g -> g = W/m. With some algebra, that means \[ g = \frac{a\cdot W}{F_{applied} - \mu_k W} \] or \[ g = \frac{a}{(F_{applied}/W) - \mu_k} \]